April 23rd 2020

  • Progress
    • VMM3
      Board powered up and board could be pinged but could not get any events. No triggers thought it was because of configuration. George gave another configuration file known to work. Issue with Virtual Machine not connecting to the board. Ben gave some suggestions to fix the issue. Data was recorded, ROOT output, need to learn how to look at the data. Histogramming not supported any more, need to use ROOT to look at data.
    • APV25
      ***No permission for Danning to access the clean room. Maybe give it a try to instruct Walt or Jack to do it.
      • right now 12 MPDs connected
    • FADC : Ed and Ben in discussion
  • NALU chip ASOC : ordered SMA adapters, will need hardware for further testing
  • Discussion VMM3 test with X rays
  • Discussion Compton asymmetries from last week
    • Question from Bryan :
    • accumulated or by pairs
    • BCM from from FADC to VTP
  • is there a change in window width ?
  • no beam run, asymmetry small
  • check if charge or position
  • pick-up from helicity
  • high frequency - signal - check asymmetry from clock ( 40 MHz in scaler )
  • small asymmetry on beam off data
  • if beam try different higher threshold
  • try to break run in several pieces
  • compute asymmetry for pairs and quad
  • helicity signal - going through different modules - could change width - but not seen in other DAQ, Ben measured MPS to be exactly 90 us so looks ok
  • Laser on /off period from script

Updated by Alexandre Camsonne over 4 years ago · 6 revisions